OpenStudy (anonymous):

Use the Pythagorean Theorem to find the missing side of the triangle. Round to the nearest tenth, if necessary. a = ?, b = 5, c = 8 A. 6.0 B. 6.1 C. 6.2 D. 6.3

OpenStudy (anonymous):

8^2=5^2+a^2

OpenStudy (anonymous):

\[Multipy and simplify: \left(\begin{matrix}8 \\x^2\end{matrix}\right) * \left(\begin{matrix}x^4 \\ 4x\end{matrix}\right) \]

OpenStudy (anonymous):

\[A. 4x^2\] \[B. 2x\] \[C. x\]

OpenStudy (anonymous):

I think it's B but im not sure

OpenStudy (anonymous):

Ugh thats no help big bro

OpenStudy (anonymous):

srry #_#

OpenStudy (anonymous):

im not good with math lol im like a 4 year old with math lol

OpenStudy (anonymous):

hahaha.... ok sis tell me what do u want me to do?

OpenStudy (anonymous):

HELP IM SOOO STUCK!!

OpenStudy (anonymous):

ok 1st one a^2+b^2=c^2 a^2+25= 64 a^2= 64-25 \[a=\sqrt{39}\]

OpenStudy (anonymous):

second one is a combination question ???

OpenStudy (anonymous):

second answer is 2x